### v1

# Modeling reflection spectra of super-Eddington X-ray sources

Abstract
We present a relativistic disk reflection model based on the geometry calculated using analytical formulae for super-Eddington accretion flows. This model features a slim disk geometry where the inner disk thickness is proportional to radius, becoming thicker as the mass accretion rate increases. The slim disk profile reduces the brightness of the blue horn in the Fe K emission line for a fixed emissivity and significantly changes the intensity profile for a lamppost geometry. The model is constructed assuming a spherically symmetric spacetime. It can be used for any kind of sources showing fluorescent reflection features and predicted to have slim accretion disks, like slow rotating black holes in X-ray binaries, active galactic nuclei, tidal disruption events, and neutron star X-ray binaries. To show the capability of the model, we use the 2017 NICER and NuSTAR data of the ultraluminous X-ray transient Swift J0243.6+6124.
Accretion (14) — Ultraluminous x-ray sources (2164) — X-ray astronomy (1810)
facilities: Swift, NICER, NuSTARsoftware: astropy (Astropy Collaboration et al., 2013, 2018), XSPEC (Arnaud, 1996), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), raytransfer (Abdikamalov et al., 2024a), blackray (Abdikamalov et al., 2024b), blacklamp (Abdikamalov et al., 2024c)
1 Introduction
Super-Eddington accretion has become an important phenomenon to be studied in the recent years in sources where the luminosity exceeds the Eddington luminosity ($L_{\rm Edd}\approx 1.26\times 10^{38}M/M_{\odot}~{}\mathrm{erg~{}s}^{-1}$). It has been observed in Ultraluminous X-ray sources (ULXs) (Kaaret et al., 2017; Fabrika et al., 2021) which are compact objects observed in nearby galaxies and even one Galactic pulsar Swift J0243.6+6124 (Cenko et al., 2017; Kennea et al., 2017). Super-Eddington accretion has been observed in tidal disruption events (TDEs), which are transient events occurring when a star is disrupted (see Gezari, 2021, for a recent review on TDEs) and accreted by a massive black hole (see Bloom et al. (2011); Burrows et al. (2011); Kara et al. (2016, 2018); Lin et al. (2015)). There are also observations of super-Eddington accretion in active galactic nuclei (AGNs) (Lanzuisi et al., 2016; Tsai et al., 2018) where the accretion takes place onto supermassive black holes at the centres of galaxies. Accretion beyond the critical Eddington limit is also considered an important factor of growth of massive black holes discovered in the early Universe (Mortlock et al., 2011; Fan et al., 2023; Massonneau et al., 2023; Matthee et al., 2024).
Fluorescent iron lines are commonly observed in the X-ray spectra of a number of astrophysical sources, such as neutron star (Bhattacharyya & Strohmayer, 2007; Cackett et al., 2008) and black hole (Tripathi et al., 2021; Draghis et al., 2023; Liu et al., 2023a) X-ray binaries (XRBs), AGNs (Nandra et al., 2007; Walton et al., 2013), and TDEs (Yao et al., 2022; Kara et al., 2016). The source of this emission line is predicted to be the “reflection” taking place at the accretion disk (or some optically thick medium). The Comptonized photons from hot plasma around the central compact object interact with the material in the disk or outflow. This imprints the atomic data of the gas as emission and absorption lines onto the spectrum. However, due to the compact object there is strong gravity which blurs this reflection spectrum arising from gravitational redshift and Doppler boosting (see, e.g., Bambi, 2017; Tanaka et al., 1995; Fabian et al., 1989). The relativistic reflection spectrum is characterized by the Compton hump and blurred elemental emission lines, out of which Fe K emission line at 6.4-6.9 keV is often the strongest feature (Ross & Fabian, 2005). Studying these reflection features is termed as X-ray reflection spectroscopy (see, e.g., Bambi et al., 2021): it provides a useful tool to probe the accretion dynamics and gravitational effects near to the compact object and remains one of the leading techniques to measure black hole spins (Brenneman & Reynolds, 2006; Draghis et al., 2023) and even performing tests of fundamental physics (Cao et al., 2018; Tripathi et al., 2019). However, the systematic effects introduced by the common assumption of a razor-thin accretion disk geometry (Novikov & Thorne, 1973; Page & Thorne, 1974) have not been thoroughly investigated. Some studies (e.g., Shashank et al. (2022); Jiang et al. (2022); Tripathi et al. (2021); Abdikamalov et al. (2020); Taylor & Reynolds (2018)) have explored this topic, but they remain within the geometrically thin regime suitable for sub-Eddington accretion disks ($H/r<<1$). Jiang et al. (2022) found that the systematic uncertainty is small compared to the statistical uncertainty of current data. Shashank et al. (2022) studied these systematic uncertainties by fitting the data with the current reflection models to spectra obtained from numerically simulated disks and found that for current missions (NuSTAR) the models work well. Effects of geometrically thick disks were explored in Riaz et al. (2020a, b). Future high-resolution data from missions like AXIS, HEX-P, or Athena will require more precise models. This is particularly important in the super-Eddington regime, as more objects such as TDEs and ULXs exhibit relativistic reflection spectra in their X-ray observations (e.g., Jaisawal et al. (2019); Yao et al. (2022, 2024); Kara et al. (2016); Motta et al. (2017); Veledina et al. (2023)). A full model which takes into account slim disks for super-Eddington scenarios have not been developed for studying reflection spectra (as per the authors’ knowledge). We note that there are indeed existing models for the thermal emissions arising from slim disk geometries (Kawaguchi, 2003; Sadowski, 2011; Caballero-Garcia et al., 2017; Wen et al., 2022). Recently, Zhang et al. (2024) studied ironline emissions from the optically thick outflows in a super-Eddington regime.
In this work, we introduce a new relativistic reflection model based on a slim disk profile for super-Eddington accretion (Watarai, 2006). The article is organised as follows. In Section 2, we describe the model REFLUX. Section 3 shows working of the model by fitting the NICER and NuSTAR data from the 2017 outburst of the galactic ULX pulsar Swift J0243.6+6124. Finally, we discuss the results and future directions in Section 4. Units used in this article are $G=c=1$.
2 Model description
To model the expected slim disk profile for a supercritical accreting source, we use the analytical disk model introduced by Watarai (2006). The model assumes an optically thick disk by introducing a radially dependent ratio $f=\mathcal{Q}^{-}_{\rm adv}/\mathcal{Q}^{+}_{\rm vis}$. $\mathcal{Q}^{-}_{\rm adv}$ is the advective cooling rate and $\mathcal{Q}^{+}_{\rm vis}$ is the viscous heating rate (for their explicit forms, see Watarai, 2006). $f$ also depends on the accretion rate, which then controls the height of the disk in our implementation. The scale height of the disk is defined as
$H=\left[(2N+3)\frac{B\Gamma_{\Omega}\Omega_{0}^{2}}{\xi}\right]^{1/2}f^{1/2}\tilde{r}\,,$ (1)
where $\Omega_{0}$ is a multiplier to the Keplerian angular velocity ($\Omega=\Omega_{0}\Omega_{\rm K}$), which we set to unity for having a disk with Keplerian velocity ($\Omega_{\rm K}$), and $\Gamma_{\Omega}=-d\log\Omega/d\log r$ is the linear approximation form of the angular velocity. We take $B=1$, $N=3$, $\xi=1.5$, and $\Gamma_{\Omega}=1.5$ as detailed in Watarai (2006). $f$ is defined as
$f(x)=\frac{1}{2}\left[D^{2}x^{2}+2-Dx(D^{2}x^{2}+4)^{1/2}\right]\,,$ (2)
where
$D=\left[\frac{64\Omega_{0}^{2}}{(2N+3)\xi B\Gamma_{\Omega}}\right]^{1/2},\quad x=\frac{\tilde{r}}{\dot{m}}\,,$ (3)
$\dot{m}=\dot{M}/\dot{M}_{\rm crit}$ is the accretion rate in units of Eddington accretion rate ($\dot{M}_{\rm crit}=L_{\rm Edd}$, where $L_{\rm Edd}$ is the Eddington luminosity of the source), and $\tilde{r}$ is the disk radius. In our implementation of the model, we only vary $\dot{m}$, which gives us a family of disks whose scale heights are controlled by the value of $\dot{m}$. Also, we keep the inner edge of the disk as a free parameter, hence the $\tilde{r}$ represented above is the disk radius starting from the inner edge. Fig. 1 shows the disk profile obtained by changing the value of the accretion rate.
To construct a full relativistic reflection model, we assume Schwarzschild spacetime around the central object. Further we use the same methods of ray-tracing (Abdikamalov et al., 2024a) and Cunningham’s transfer function (Cunningham, 1975; Speith et al., 1995; Dauser et al., 2010) as demonstrated in detail in Bambi et al. (2017); Abdikamalov et al. (2019, 2020) (specifically, exactly the same methods have been used to calculate the transfer functions as discussed in Sec. 3 of Abdikamalov et al. (2020) for a thin disk with finite thickness).
The free parameters of the model are: the inner edge of the disk, which in the current version of REFLUX varies from $6M$ to $500M$, the accretion rate, which is varied from $1\dot{M}_{\rm crit}$ to $100\dot{M}_{\rm crit}$, and the inclination of the disk, varying from $3^{\circ}$ to $70^{\circ}$. While constructing the model, we also have the outer radius of the disk up to $1000M$. For the lamppost versions, we have the height of the lamppost corona ranging from $2.2M$ to $100M$. With higher values of inclination, the photons from the inner regions of the disk do not reach the distant observer, so we do not consider inclinations higher than $70^{\circ}$. In Fig. 2, we show the variations in iron line profiles for different parameter values, we see that the thickness of the disk produces a change in the shape of the ironlines, reducing the blue horn from thinner to a thicker disk. Fig. 3 shows the ironline profiles for a lamppost emissivity. In the lamppost corona geometry, the emissivity profile of the disk significantly changes as inner region of the disk is more illuminated due to the thicker disk. For low corona heights and high accretion rates, the outer regions of the disk are shadowed due to the thickness of the disk.
REFLUX works along with the xillver (García et al., 2013) tables for calculating the reflection spectra in the rest frame of the gas. Tab. 1 guides the full set of parameters available across the different sub-models (flavors) of REFLUX  [^1] With the exception of $\dot{m}$ which is the accretion rate, the parameters listed in Tab. 1 are same as that of the relxill model listed here: https://www.sternwarte.uni-erlangen.de/~dauser/research/relxill/index.html.. Each of the sub-model mainly arises from using different xillver tables. Sub-models with LP suffix mean they use a lamp-post emissivity. REFLUX is the base model using the angle-resolved cutoff powerlaw xillver table. REFLUXCp uses the xillverCp table with nthComp (Zdziarski et al., 1996; Życki et al., 1999) Comptonization for incident spectrum. REFLUXD uses the xillverD table, which has electron density as a free parameter and cutoff power-law with energy cutoff of 300 keV. RELLINE_SD is the relativistic line model.
Figure 1: Disk profile obtained from the model for various accretion rates. The accretion rate $\dot{m}$ is in units of $\dot{M}_{\rm crit}$. The inner edge is set to $10M$. The black region at the origin denotes the event horizon of a Schwarzschild black hole.
Figure 2: Iron lines obtained from the model for different parameter values. The left panel shows some iron line profiles for different values of the mass accretion rate, the central panel shows some iron line profiles for different values of inner edge of the disk, and the right panel shows some iron line profiles for different values of the disk inclination angle. The other fixed parameter values are on top of every panel. The emissivity index is set to $q_{in}=q_{out}=3$ in all cases.
Figure 3: Iron lines obtained from the model for different disk heights using the lamppost emissivity. The left panel shows some iron line profiles for corona height $h=6M$, the central panel shows iron line profiles for $h=10M$, and the right panel shows iron line profiles $h=50M$. The other fixed parameter values are $\iota=45^{\circ}$, $R_{in}=10M$ and $\Gamma=2$
[…]
RELLINE_SD $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$ $\times$
REFLUX $\times$ $\times$ 15 $\times$ $\times$
REFLUXCp $\times$ $\times$ 15 $\times$ $\times$
REFLUXD $\times$ $\times$ 300 $\times$ $\times$
RELLINELP_SD $\times$ $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$ $\times$
REFLUXLP $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$
REFLUXLPCp $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$
REFLUXLPD $\times$ $\times$ $\times$ $\times$ 300 $\times$ $\times$
Table 1: Parameter availability for different sub-models. Here, ✓means the parameter is available in the given sub-model and $\times$ means the parameter is not available.
3 Spectral fitting of a ULX pulsar
In this section, we discuss the results of our fitting we obtained for the Galactic ULX pulsar Swift J0243.6+6124 (Cenko et al., 2017; Kennea et al., 2017). The source went into a bright outburst in 2017, with its peak luminosity $>10^{39}$ erg s-1, far exceeding the Eddington limit for a neutron star (Wilson-Hodge et al., 2018).
We analyze the quasi-simultaneous NICER (ObsID 1050390113, 1 November 2017) and NuSTAR (ObsID 90302319004, 30 October 2017) observations that have been fitted with the standard thin-disk reflection model in Jaisawal et al. (2019) (see their Fig. 7). We follow the procedure in Jaisawal et al. (2019) to process the NuSTAR (FPMA and FPMB) data. The NICER data are first processed with nicerl2 (CALDB v20240206 and geomagnetic data updated to March 2024). The source and background spectral files are then extracted using nicel3-spect with the default background model. Spectra are grouped to ensure a minimal counts of 100 for each bin before fitting with XSPEC v12.13.0c (Arnaud, 1996). Best-fit parameter values and uncertainties (90% confidence level unless otherwise stated) are obtained using $\chi^{2}$ statistics.
As in Jaisawal et al. (2019), we first fit the NICER (1–10 keV) and NuSTAR (3–79 keV) data with the model: constant * tbabs * (bbodyrad + cutoffpl + relxill/relxilllp). In this model, the relxill and relxilllp (García et al., 2014) component represent relativistic reflection from a razor-thin accretion disk, assuming fixed emissivity ($q=3$) and the corona as a point source above the black hole at a certain height ($h$), respectively. We fix the spin parameter at $a_{*}=0$ and the outer radius of the accretion disk at 1000 $M$. The best-fit parameters and the corresponding uncertainties are shown in Tab. 2 for fixed emissivity and in Tab. 3 with lamp-post corona geometry.
At the time of this observation, the mass accretion rate of the source was at super-Eddington level. The absorption corrected X-ray flux in the 0.01-100 keV band is around 10 times of the Eddington limit. Therefore, a thick disk, instead of thin one, is expected. We replace the relxill/relxilllp component with our REFLUX/REFLUXLP model, with the mass accretion rate fixed at 10 times of the Eddington accretion rate. As shown in Tabs. 2 and 3, the models provide a good fit to the data, with the $\chi^{2}$ close to that of relxill/relxilllp. The best-fit model components and residuals are shown in Fig. 4.
[…]
Tbabs nH $10^{22}$ cm-2 $0.683_{-0.011}^{+0.009}$ $0.664_{-0.006}^{+0.01}$
Cutoffpl $\Gamma$ $1.423_{-0.009}^{+0.009}$ $1.434_{-0.009}^{+0.014}$
Ecut keV $24.3_{-0.5}^{+0.6}$ $24.8_{-0.4}^{+0.3}$
norm $12.4_{-0.5}^{+0.5}$ $12.40_{-0.29}^{+0.5}$
Bbodyrad kT keV $1.213_{-0.018}^{+0.019}$ $1.234_{-0.013}^{+0.009}$
norm $955_{-33}^{+21}$ $988_{-27}^{+25}$
Reflection Incl deg $15.1_{-2.4}^{+4}$ $25.5_{-1.7}^{+3}$
$R_{\rm in}$ $M$ $37_{-7}^{+9}$ $34_{-7}^{+8}$
$\log(\xi)$ erg cm s-1 $3.54_{-0.07}^{+0.05}$ $3.29_{-0.12}^{+0.03}$
$A_{\rm Fe}$ solar $4.2_{-1.2}^{+1.4}$ $2.39_{-0.27}^{+0.5}$
norm $0.050_{-0.005}^{+0.007}$ $0.0255_{-0.0017}^{+0.001}$
$C_{\rm FPMB}$ $0.9880_{-0.0014}^{+0.0014}$ $0.9880_{-0.0014}^{+0.0014}$
$C_{\rm XTI}$ $1.242_{-0.012}^{+0.012}$ $1.240_{-0.012}^{+0.01}$
$\chi^{2}/\nu$ 2855.14/2624 2857.12/2624
Table 2: Best-fit parameters for the model with relxill and the model with REFLUX. The reported uncertainties correspond to the 90% interval limits for one relevant parameter ($\Delta\chi^{2}=2.71$).
[…]
Tbabs nH $10^{22}$ cm-2 $0.667_{-0.009}^{+0.009}$ $0.671_{-0.008}^{+0.007}$
Cutoffpl $\Gamma$ $1.434_{-0.009}^{+0.01}$ $1.438_{-0.01}^{+0.008}$
Ecut keV $24.74_{-0.4}^{+0.26}$ $24.7_{-0.3}^{+0.3}$
norm $12.4_{-0.3}^{+0.4}$ $12.6_{-0.3}^{+0.3}$
Bbodyrad kT keV $1.23_{-0.01}^{+0.01}$ $1.235_{-0.006}^{+0.005}$
norm $1000_{-30}^{+40}$ $970_{-25}^{+30}$
Reflection $h$ $M$ $2.20_{-P}^{+0.07}$ $>65$
Incl deg $25.7_{-1.3}^{+2.7}$ $26.5_{-2.4}^{+1.2}$
$R_{\rm in}$ $M$ $39_{-8}^{+9}$ $<14$
$\log(\xi)$ erg cm s-1 $3.31_{-0.05}^{+0.04}$ $3.284_{-0.07}^{+0.029}$
$A_{\rm Fe}$ solar $2.4_{-0.3}^{+0.3}$ $2.6_{-0.3}^{+0.4}$
norm $3.58_{-0.7}^{+0.24}$ $1.37_{-0.06}^{+0.10}\times 10^{-7}$
$C_{\rm FPMB}$ $0.9880_{-0.0014}^{+0.0014}$ $0.9880_{-0.0014}^{+0.0014}$
$C_{\rm XTI}$ $1.223_{-0.013}^{+0.012}$ $1.238_{-0.01}^{+0.011}$
$\chi^{2}/\nu$ 2816.49/2623 2826.35/2623
Table 3: Best-fit parameters for the model with relxilllp and the model with REFLUXLP. The reported uncertainties correspond to the 90% interval limits for one relevant parameter ($\Delta\chi^{2}=2.71$).
Figure 4: Best-fit model components and residuals for the lamppost geometry (left) and the power-law emissivity profile (right) with the REFLUX model. The black and red colors represent data of FPMA and FPMB, respectively. NICER data are plotted in cyan. Data are rebinned for visual clarity.
4 Discussions
The quality of the fits with the model with relxill and the model with REFLUX is similar. We fit the same data and the two models have the same number of free parameters, so we can directly compare their $\chi^{2}$ and the difference is marginal ($\Delta\chi^{2}=1.98$ for fixed emissivity and $9.86$ for lamppost). The $\chi^{2}$ of the best-fit with relxill is slightly lower – a possible reason for this could be the smaller parameter grid of REFLUX than relxill in the FITS file. However, we know that the source was accreting around 10 times of its Eddington limit, so the fit with REFLUX appears to be more appropriate for the spectra analyzed. In Tab. 2, the two models provide consistent estimates of the parameters with the exception of the inclination angle of the disk and the iron abundance. The measurements of the inclination angle are $i=15.1_{-2.4}^{+4}$ deg (relxill) and $25.5_{-1.7}^{+3}$ (REFLUX). However, the estimates of disk inclination angles inferred from the analysis of reflection features are normally not very accurate and both models suggest a low value of $i$. Estimates of the iron abundance using X-ray reflection spectroscopy are even more problematic. It is currently unclear why reflection measurements often suggest super-solar iron abundances (but there are also sources for which reflection measurements suggest iron abundances significantly lower than the solar one); see, for instance, Bambi et al. (2021) for a short review on the problem. It may be related to a deficiency of current reflection models (e.g., underestimate of the disk electron density (Jiang et al., 2019a, b; Liu et al., 2023b), presence of two or more coronae (Fürst et al., 2015), etc.) or a true physical phenomenon, like radiative levitation of metal ions in the inner part of the accretion disk (Reynolds et al., 2012). The estimate of the iron abundance with REFLUX would go to the right direction, decreasing the value of $A_{\rm Fe}$ with respect to that inferred with relxill, but we would need to analyze more sources and more observations to see if this is a feature of our model and not an accidental result for the spectra analyzed in the previous section. In Tab. 3, we see that for a lamppost geometry, the inclination angle and the iron abundance for both the models closely match the result of REFLUX in the analysis with fixed emissivity. However, we see a significant difference in the corona height and the inner edge of the disk. With relxill, we see a very low corona height (the lowest in the model) and a larger value for the inner edge. With REFLUX, the case is the opposite, we obtain a higher corona height $h>65M$ and inner edge closer to the central object $R_{in}<14M$. In the case of REFLUX, we see a small norm possibly because the whole disk may not be illuminated for many different $h$ and $R_{in}$ configurations making the flux from the outer part of the disk negligible. We caution the readers against interpreting the high value of $h>65M$ too rigidly. As in our analysis, we assume a lamppost geometry, which may not be applicable to an Ultra-Luminous Pulsar (ULP) like Swift J0243.6+6124. The system might instead contain a more extended primary illuminating source, such as an accretion column (e.g., Pinto & Walton (2023)). Nonetheless, we observed changes in inferred geometries when a slim disk model was employed. This underscores the critical importance of considering disk geometry when addressing systematic uncertainties. Such uncertainties significantly affect the inferred disk inner radius and, consequently, the magnetic field of ULPs (e.g., Jaisawal et al. (2019)). We will further investigate Swift J0243.6+6124 in this perspective in the future.
A more detailed analysis of the 2017 NICER and NuSTAR data of Swift J0243.6+6124 is beyond the purpose of the present work, which is the presentation of REFLUX. In our fits (with both relxill and REFLUX), we have assumed that the spectrum illuminating the disk, and responsible for the reflection component can be described by a power law with a high-energy cutoff. However, this may not be the case for Swift J0243.6+6124. A power law with a high-energy cutoff is the spectrum expected for Comptonized photons, which is the case in which thermal photons from an accretion disk can inverse Compton scatter off free electrons in a hot corona. In the case of a neutron star, reflection features may be produced by illumination of the disk by thermal photons from the surface of the neutron star. Such a scenario could be implemented in REFLUX by using the xillverNS table, in which reflection spectra are produced by an incident blackbody spectrum.
While for neutron star XRBs like Swift J0243.6+6124 the Schwarzschild metric should describe well the spacetime geometry around the source, in the case of other super-Eddington X-ray sources (black hole XRBs, AGNs, and TDEs) deviations from the Schwarzschild background can be important and it will be necessary to implement the Kerr metric in REFLUX. This is also left to future work.
This work was supported by the National Natural Science Foundation of China (NSFC), Grant No. 12250610185 and 12261131497, and the Natural Science Foundation of Shanghai, Grant No. 22ZR1403400. J.J. acknowledges support from Leverhulme Trust, Isaac Newton Trust and St Edmund’s College, University of Cambridge.

### v2

@@
 Abstract
-We present a relativistic disk reflection model based on the geometry calculated using analytical formulae for super-Eddington accretion flows.
-This model features a slim disk geometry where the inner disk thickness is proportional to radius, becoming thicker as the mass accretion rate increases.
-The slim disk profile reduces the brightness of the blue horn in the Fe K emission line for a fixed emissivity and significantly changes the intensity profile for a lamppost geometry.
-The model is constructed assuming a spherically symmetric spacetime.
-It can be used for any kind of sources showing fluorescent reflection features and predicted to have slim accretion disks, like slow rotating black holes in X-ray binaries, active galactic nuclei, tidal disruption events, and neutron star X-ray binaries.
-To show the capability of the model, we use the 2017 NICER and NuSTAR data of the ultraluminous X-ray transient Swift J0243.6+6124.
+Relativistic reflection is a common feature in the X-ray observations of accreting compact objects.
+We present reflux, a new X-ray reflection model for spectral analysis of super-Eddington sources.
+We develop two relativistic reflection frameworks for super-Eddington accretion: a slim-disk model that self-consistently accounts for disk thickening and self-shadowing, and an optically thick wind model that treats reflection off a funnel-shaped outflow.
+The slim-disk model offers a geometry where the inner disk thickness is proportional to radius, becoming thicker as the mass accretion rate increases.
+The wind model measures the opening angle of the funnel, the wind speed, and wind acceleration radius.
+The slim-disk profile reduces the brightness of the blue horn in the Fe K emission line for a fixed emissivity and significantly changes the intensity profile for a lamppost geometry.
+The wind model shows a blue-shifted iron line due to high velocity outflows.
+Both models assume a spherically symmetric spacetime.
+We apply the wind model to the XMM-Newton spectrum of the tidal disruption event Swift J1644+57, where the Fe K profile is expected to be shaped by scattering in an outflowing funnel.
+We constrain the opening angle of the funnel and find a high velocity of the wind.
 Accretion (14) — Ultraluminous x-ray sources (2164) — X-ray astronomy (1810)
-facilities: Swift, NICER, NuSTARsoftware: astropy (Astropy Collaboration et al., 2013, 2018), XSPEC (Arnaud, 1996), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), raytransfer (Abdikamalov et al., 2024a), blackray (Abdikamalov et al., 2024b), blacklamp (Abdikamalov et al., 2024c)
+facilities: Swift, NICER, NuSTAR, XMM-Newtonsoftware: astropy (Astropy Collaboration et al., 2013, 2018), XSPEC (Arnaud, 1996), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), raytransfer (Abdikamalov et al., 2024a), blackray (Abdikamalov et al., 2024b), blacklamp (Abdikamalov et al., 2024c), ziji (Mirzaev et al., 2024a)
 1 Introduction
@@
 It has been observed in Ultraluminous X-ray sources (ULXs) (Kaaret et al., 2017; Fabrika et al., 2021) which are compact objects observed in nearby galaxies and even one Galactic pulsar Swift J0243.6+6124 (Cenko et al., 2017; Kennea et al., 2017).
-Super-Eddington accretion has been observed in tidal disruption events (TDEs), which are transient events occurring when a star is disrupted (see Gezari, 2021, for a recent review on TDEs) and accreted by a massive black hole (see Bloom et al.
-(2011); Burrows et al.
-(2011); Kara et al.
-(2016, 2018); Lin et al.
-(2015)).
-There are also observations of super-Eddington accretion in active galactic nuclei (AGNs) (Lanzuisi et al., 2016; Tsai et al., 2018) where the accretion takes place onto supermassive black holes at the centres of galaxies.
+Super-Eddington accretion has been observed in tidal disruption events (TDEs), which are transient events occurring when a star is disrupted (see Gezari, 2021, for a recent review on TDEs) and accreted by a massive black hole (see Bloom et al., 2011; Burrows et al., 2011; Kara et al., 2016, 2018; Lin et al., 2015).
+There are also observations of super-Eddington accretion in active galactic nuclei (AGNs) (Lanzuisi et al., 2016; Tsai et al., 2018) where accretion takes place onto supermassive black holes at the centers of galaxies.
 Accretion beyond the critical Eddington limit is also considered an important factor of growth of massive black holes discovered in the early Universe (Mortlock et al., 2011; Fan et al., 2023; Massonneau et al., 2023; Matthee et al., 2024).
-Fluorescent iron lines are commonly observed in the X-ray spectra of a number of astrophysical sources, such as neutron star (Bhattacharyya & Strohmayer, 2007; Cackett et al., 2008) and black hole (Tripathi et al., 2021; Draghis et al., 2023; Liu et al., 2023a) X-ray binaries (XRBs), AGNs (Nandra et al., 2007; Walton et al., 2013), and TDEs (Yao et al., 2022; Kara et al., 2016).
-The source of this emission line is predicted to be the “reflection” taking place at the accretion disk (or some optically thick medium).
+Fluorescent iron lines are commonly observed in the X-ray spectra of a number of astrophysical sources, such as neutron star (Bhattacharyya & Strohmayer, 2007; Cackett et al., 2008) and black hole (Tripathi et al., 2021b; Draghis et al., 2023; Liu et al., 2023a, b) X-ray binaries (XRBs), AGNs (Nandra et al., 2007; Walton et al., 2013), and TDEs (Yao et al., 2022; Kara et al., 2016).
+The source of this emission line is predicted to be the “reflection” taking place at the accretion disk (or some optically thick medium like winds).
 The Comptonized photons from hot plasma around the central compact object interact with the material in the disk or outflow.
@@
 The relativistic reflection spectrum is characterized by the Compton hump and blurred elemental emission lines, out of which Fe K emission line at 6.4-6.9 keV is often the strongest feature (Ross & Fabian, 2005).
-Studying these reflection features is termed as X-ray reflection spectroscopy (see, e.g., Bambi et al., 2021): it provides a useful tool to probe the accretion dynamics and gravitational effects near to the compact object and remains one of the leading techniques to measure black hole spins (Brenneman & Reynolds, 2006; Draghis et al., 2023) and even performing tests of fundamental physics (Cao et al., 2018; Tripathi et al., 2019).
+Studying these reflection features is termed as X-ray reflection spectroscopy (see, e.g., Bambi et al., 2021): it provides a useful tool to probe the accretion dynamics and gravitational effects near to the compact object and remains one of the leading techniques to measure black hole spins (Brenneman & Reynolds, 2006; Draghis et al., 2023) and even performing tests of fundamental physics (Cao et al., 2018; Tripathi et al., 2019b, a).
+Current reflection models are designed to work in the thin-disk regime, which is normally thought to hold up to 20-30% of the Eddington accretion rate (Shakura & Sunyaev, 1973; Novikov & Thorne, 1973).
+At higher accretion rates, the assumptions of the standard thin‐disk model begin to break down.
+When a source exceeds its Eddington limit, radiative diffusion becomes inefficient and a non‐negligible fraction of the dissipated energy is advected inward rather than being locally radiated (Abramowicz et al., 1988).
+In this super-Eddington regime, the accretion flow will instead adopt a slim-disk structure and the disk no longer remains geometrically thin, in which radial advection and pressure gradients play a dominant role (Abramowicz et al., 1988; Watarai, 2006; Sadowski, 2009).
+More recently, from the studies using radiative general relativistic magnetohydrodynamic simulations of super-Eddington accretion flows, it is also predicted that fast and optically think winds can be launched from the super-Eddington system due to magnetic and radiation pressures (McKinney et al., 2014; Jiang et al., 2014, 2019; Sadowski et al., 2015; Sadowski & Narayan, 2016; Thomsen et al., 2022).
+This optically thick wind creates a funnel-like geometry in which reflection happens (see e.g.
+Fig.
+4).
 However, the systematic effects introduced by the common assumption of a razor-thin accretion disk geometry (Novikov & Thorne, 1973; Page & Thorne, 1974) have not been thoroughly investigated.
-Some studies (e.g., Shashank et al.
-(2022); Jiang et al.
-(2022); Tripathi et al.
-(2021); Abdikamalov et al.
-(2020); Taylor & Reynolds (2018)) have explored this topic, but they remain within the geometrically thin regime suitable for sub-Eddington accretion disks ($H/r<<1$).
+Some studies (e.g., Shashank et al., 2022; Jiang et al., 2022; Tripathi et al., 2021a; Abdikamalov et al., 2020; Taylor & Reynolds, 2018) have explored this topic, but they remain within the geometrically thin regime suitable for sub-Eddington accretion disks ($H/r<<1$).
 Jiang et al.
@@
 Shashank et al.
-(2022) studied these systematic uncertainties by fitting the data with the current reflection models to spectra obtained from numerically simulated disks and found that for current missions (NuSTAR) the models work well.
+(2022, 2025) studied these systematic uncertainties by fitting the data with the current reflection models to spectra obtained from numerically simulated disks and found that for current missions (NuSTAR) the models work well for fast spinning black holes.
 Effects of geometrically thick disks were explored in Riaz et al.
 (2020a, b).
-Future high-resolution data from missions like AXIS, HEX-P, or Athena will require more precise models.
-This is particularly important in the super-Eddington regime, as more objects such as TDEs and ULXs exhibit relativistic reflection spectra in their X-ray observations (e.g., Jaisawal et al.
-(2019); Yao et al.
-(2022, 2024); Kara et al.
-(2016); Motta et al.
-(2017); Veledina et al.
-(2023)).
-A full model which takes into account slim disks for super-Eddington scenarios have not been developed for studying reflection spectra (as per the authors’ knowledge).
-We note that there are indeed existing models for the thermal emissions arising from slim disk geometries (Kawaguchi, 2003; Sadowski, 2011; Caballero-Garcia et al., 2017; Wen et al., 2022).
+Future high-resolution data from missions like eXTP, AXIS, or NewAthena will require more precise models.
+This is particularly important in the super-Eddington regime, as more objects such as TDEs and ULXs exhibit relativistic reflection spectra in their X-ray observations (e.g., Jaisawal et al., 2019; Yao et al., 2022, 2024; Kara et al., 2016; Motta et al., 2017; Veledina et al., 2024).
+A full model which takes into account slim disks and optically thick winds for super-Eddington scenarios have not been developed for studying reflection spectra (as per the authors’ knowledge).
+We note that there are indeed existing models for the thermal emissions arising from slim-disk geometries (Kawaguchi, 2003; Sadowski, 2011; Caballero-Garcia et al., 2017; Wen et al., 2022).
 Recently, Zhang et al.
-(2024) studied ironline emissions from the optically thick outflows in a super-Eddington regime.
-In this work, we introduce a new relativistic reflection model based on a slim disk profile for super-Eddington accretion (Watarai, 2006).
-The article is organised as follows.
+(2024) studied iron line emissions modeled from the optically thick winds in a super-Eddington regime.
+In this work, we introduce reflux, a new relativistic reflection model with two geometries: $(i)$ slim-disk profile for super-Eddington accretion following Watarai (2006) and $(ii)$ optically thick wind model similar to Zhang et al.
+(2024).
+The article is organized as follows.
 In Section 2, we describe the model reflux.
-Section 3 shows working of the model by fitting the NICER and NuSTAR data from the 2017 outburst of the galactic ULX pulsar Swift J0243.6+6124.
+Section 3 shows working of the model by fitting the 2011 XMM-Newton observation of TDE Swift J1644+57.
 Finally, we discuss the results and future directions in Section 4.
-Units used in this article are $G=c=1$.
+Units used in this article are $G=c=1$ unless mentioned otherwise.
 2 Model description
-To model the expected slim disk profile for a supercritical accreting source, we use the analytical disk model introduced by Watarai (2006).
+To construct a full relativistic reflection model, we assume Schwarzschild spacetime around the central object.
+Further we use the same methods of ray-tracing (Abdikamalov et al., 2024a) and Cunningham’s transfer function (Cunningham, 1975; Speith et al., 1995; Dauser et al., 2010) as demonstrated in detail in Bambi et al.
+(2017); Abdikamalov et al.
+(2019, 2020) (specifically, exactly the same methods have been used to calculate the transfer functions as discussed in Sec.
+3 of Abdikamalov et al., 2020).
+The new parameters in reflux related to the super-Eddington accretion geometries are discussed in the next sections.
+The parameters related to xillver tables are not discussed here as they are same as relxill [^1] The parameters of the relxill and xillver model can be found here: https://www.sternwarte.uni-erlangen.de/~dauser/research/relxill/index.html..
+2.1 Slim-disk
+To model the expected slim-disk profile for a supercritical accreting source, we use the analytical disk model introduced by Watarai (2006).
 The model assumes an optically thick disk by introducing a radially dependent ratio $f=\mathcal{Q}^{-}_{\rm adv}/\mathcal{Q}^{+}_{\rm vis}$.
@@
 1 shows the disk profile obtained by changing the value of the accretion rate.
-To construct a full relativistic reflection model, we assume Schwarzschild spacetime around the central object.
-Further we use the same methods of ray-tracing (Abdikamalov et al., 2024a) and Cunningham’s transfer function (Cunningham, 1975; Speith et al., 1995; Dauser et al., 2010) as demonstrated in detail in Bambi et al.
-(2017); Abdikamalov et al.
-(2019, 2020) (specifically, exactly the same methods have been used to calculate the transfer functions as discussed in Sec.
-3 of Abdikamalov et al.
-(2020) for a thin disk with finite thickness).
-The free parameters of the model are: the inner edge of the disk, which in the current version of REFLUX varies from $6M$ to $500M$, the accretion rate, which is varied from $1\dot{M}_{\rm crit}$ to $100\dot{M}_{\rm crit}$, and the inclination of the disk, varying from $3^{\circ}$ to $70^{\circ}$.
-While constructing the model, we also have the outer radius of the disk up to $1000M$.
-For the lamppost versions, we have the height of the lamppost corona ranging from $2.2M$ to $100M$.
-With higher values of inclination, the photons from the inner regions of the disk do not reach the distant observer, so we do not consider inclinations higher than $70^{\circ}$.
-In Fig.
-2, we show the variations in iron line profiles for different parameter values, we see that the thickness of the disk produces a change in the shape of the ironlines, reducing the blue horn from thinner to a thicker disk.
-Fig.
-3 shows the ironline profiles for a lamppost emissivity.
-In the lamppost corona geometry, the emissivity profile of the disk significantly changes as inner region of the disk is more illuminated due to the thicker disk.
-For low corona heights and high accretion rates, the outer regions of the disk are shadowed due to the thickness of the disk.
-REFLUX works along with the xillver (García et al., 2013) tables for calculating the reflection spectra in the rest frame of the gas.
-Tab.
-1 guides the full set of parameters available across the different sub-models (flavors) of REFLUX  [^1] With the exception of $\dot{m}$ which is the accretion rate, the parameters listed in Tab.
-1 are same as that of the relxill model listed here: https://www.sternwarte.uni-erlangen.de/~dauser/research/relxill/index.html..
-Each of the sub-model mainly arises from using different xillver tables.
-Sub-models with LP suffix mean they use a lamp-post emissivity.
-REFLUX is the base model using the angle-resolved cutoff powerlaw xillver table.
-REFLUXCp uses the xillverCp table with nthComp (Zdziarski et al., 1996; Życki et al., 1999) Comptonization for incident spectrum.
-REFLUXD uses the xillverD table, which has electron density as a free parameter and cutoff power-law with energy cutoff of 300 keV.
-RELLINE_SD is the relativistic line model.
 Figure 1: Disk profile obtained from the model for various accretion rates.
@@
 The black region at the origin denotes the event horizon of a Schwarzschild black hole.
+The free parameters of the model are: the inner edge of the disk ($R_{\rm in}$), which in the current version of reflux varies from $6M$ to $500M$, the accretion rate ($\dot{m}$), which is varied from $1\dot{M}_{\rm crit}$ to $100\dot{M}_{\rm crit}$, and the inclination of the disk ($i$), varying from $3^{\circ}$ to $70^{\circ}$.
+While constructing the model, we also have the outer radius of the disk ($R_{\rm out}$) up to $1000M$.
+For the lamppost version, we have the height of the lamppost corona ($h$) ranging from $2.2M$ to $100M$.
+With higher values of inclination, the photons from the inner regions of the disk do not reach the distant observer, so we do not consider inclinations higher than $70^{\circ}$.
+In Fig.
+2, we show the variations in iron line profiles for different parameter values: we see that the thickness of the disk produces a change in the shape of the iron lines, reducing the blue horn from thinner to a thicker disk.
+Fig.
+3 shows the iron line profiles for a lamppost emissivity.
+In the lamppost corona geometry, the emissivity profile of the disk significantly changes as the inner region of the disk is more illuminated due to the thicker disk.
+For low corona heights and high accretion rates, the outer regions of the disk are shadowed due to the thickness of the disk.
 Figure 2: Iron lines obtained from the model for different parameter values.
@@
 Figure 3: Iron lines obtained from the model for different disk heights using the lamppost emissivity.
-The left panel shows some iron line profiles for corona height $h=6M$, the central panel shows iron line profiles for $h=10M$, and the right panel shows iron line profiles $h=50M$.
-The other fixed parameter values are $\iota=45^{\circ}$, $R_{in}=10M$ and $\Gamma=2$
+The left panel shows some iron line profiles for corona height $h=6M$, the central panel shows iron line profiles for $h=10M$, and the right panel shows iron line profiles for $h=50M$.
+The other fixed parameter values are $i=45^{\circ}$, $R_{in}=10M$ and $\Gamma=2$
+2.2 Optically thick wind
+To model the optically thick winds from a super-Eddington accreting system, we follow Zhang et al.
+(2024).
+The winds launched from the system form a funnel structure within which the reflection takes place.
+Fig.
+4 shows the geometry of the wind used to model the optically thick winds which produce reflection features.
+A hot and compact corona produces the power-law photons (blue arrows) which then reflect from the winds (green arrows), then these reflected photons return back to the reflection surface and undergo secondary reflection (yellow arrows).
+The funnel follows the velocity profile of the extended CAK law (Castor et al., 1975; Thomsen et al., 2019; Parkinson et al., 2022)
+$v=v_{0}+(v_{\infty}-v_{0})\frac{[(r-r_{\rm ISCO})/R_{\rm acc}]^{\alpha}}{[(r-r_{\rm ISCO})/R_{\rm acc}]^{\alpha}+1}\,$, (4)
+where $v_{0}$ is the wind launch velocity, which we take as 0, and $v_{\infty}$ is the terminal velocity.
+$R_{\rm acc}$ is the wind acceleration radius.
+$r_{\rm ISCO}$ is the radius of the innermost stable circular orbit, which, in our case, is set to $6M$.
+The model has two emissivity profiles: broken power-law and lamppost.
+Figure 4: Sketch depicting the geometry of the optically thick wind used for reflection modeling.
+As discussed in Zhang et al.
+(2024), the secondary reflection becomes important in a wind like geometry (Fig.
+4, yellow arrows), hence, in our lamppost models we use ziji (Mirzaev et al., 2024c, b) to calculate the correct emissivity profiles for returning radiation and implement it in the model.
+This is similar to the returning radiation in the lamppost models of relxill (Dauser et al., 2022).
+The free parameters and their range in the model are follows.
+The half of opening angle of the wind funnel $\theta_{\rm half}$ is in the range $18^{\circ}$ to $45^{\circ}$.
+Acceleration radius ($R_{\rm acc}$) varies from $2M$ to $20M$.
+Terminal wind speed ($v_{\infty}$) is in the range 0 to $0.5c$.
+The inclination angle ($i$) has lowest value $3^{\circ}$ and maximum is always equal to $\theta_{\rm half}$.
+In the lamppost model the height of the corona $h$ varies from $4M$ to $20M$.
+We also have the outer radius ($R_{\rm out}$) of the accretion flow up to $400M$.
+Fig.
+5 shows the dependence of the iron lines on different parameter values in the model.
+Figure 5: Iron line profiles for the wind reflection with a lamppost.
+Each panel varies one parameter while others are fixed.
+Top-left: dependence on lamppost height ($h$).
+Top-middle: dependence on wind opening angle ($\theta_{\rm half}$).
+Top-right: dependence on acceleration radius ($R_{\rm acc}$).
+Bottom-left: dependence on terminal wind speed ($v_{\infty}$).
+Bottom-right: dependence on inclination angle (ii).
+The parameters are fixed as follows: $\Gamma=2$, $h=6M$, $\theta_{\rm half}=30^{\circ}$, $R_{\rm acc}=10M$, $v_{\infty}=0.1c$, and $i=20^{\circ}$.
+3 Super-Eddington TDE Swift J1644+57
+In this section, we discuss the fitting of for TDE Swift J1644+57 which was first discovered on 28 March 2011 by Swift Burst Alert Telescope (Burrows et al., 2011) at the center of a compact, non-interacting, star-forming galaxy with a redshift of $z=0.3534$.
+The source reached a peak luminosity of $\sim 10^{48}$ erg s-1.
+Follow up studies were done by Suzaku and XMM-Newton instruments (Reis et al., 2012).
+Here, we re-analyze the XMM-Newton observation (ObsID 0678380101, 16 March 2011).
+The XMM-Newton data were reduced using the Science Analysis System (SAS version 19.1.0) and the calibration files as of 2023.
+For the EPIC-pn camera, calibrated event files were produced with the EPPROC task.
+Background flaring was filtered by applying a threshold cut on the 10–12 keV light curve.
+The event list was then filtered using standard criteria PATTERN <= 4 and FLAG == 0.
+The source spectrum was extracted from an annular region with an outer radius of 60 arcseconds, excluding a central circular region of 12 arcseconds to mitigate pile-up.
+The background spectrum was extracted from a circular region of the same size (60 arcseconds) located near the source.
+The redistribution matrix file (RMF) and ancillary response file (ARF) were generated using the RMFGEN and ARFGEN tasks, respectively.
+We first fit the data with TBabs*zTBabs*cutoffpl with redshift set to 0.3534.
+We use TBabs to take into account galactic absorption and fix the value of $N_{H}$ to 0.0173222https://www.swift.ac.uk/analysis/nhtot/index.php.
+Fig.
+6 shows the data and the iron line obtained $\sim 8$ keV, as also noted in Kara et al.
+(2016).
+It was speculated from the energy and reverberation spectra that the source has a wind-funnel geometry (Kara et al., 2016), hence a good choice for fitting using the reflux_wind model.
+To fit the reflection/iron line component, we use the model: TBabs*zTBabs*(zcutoffpl+reflection).
+We perform three different fits with reflection set as zGauss for fitting only the iron line, relxilllp for testing with the thin-disk geometry and refluxlp_wind to fit the wind funnel geometry.
+The best-fit results are shown in Tab.
+1 and Fig.
+7.
+Figure 6: Swift J1644+57 data fitted with TBabs*zTBabs*zcutoffpl.
+Both panels show the same data; the right panel is zoomed in 5-10 keV.
+The red dashed boxes on the left panel denote the zoomed region.
+Figure 7: Best-fit model components and residuals for Swift J1644+57 fitted with reflux_wind (left), relxilllp (middle) and zGauss (right).
 […]
-RELLINE_SD $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$ $\times$
-REFLUX $\times$ $\times$ 15 $\times$ $\times$
-REFLUXCp $\times$ $\times$ 15 $\times$ $\times$
-REFLUXD $\times$ $\times$ 300 $\times$ $\times$
-RELLINELP_SD $\times$ $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$ $\times$
-REFLUXLP $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$
-REFLUXLPCp $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$
-REFLUXLPD $\times$ $\times$ $\times$ $\times$ 300 $\times$ $\times$
-Table 1: Parameter availability for different sub-models. Here, ✓means the parameter is available in the given sub-model and $\times$ means the parameter is not available.
-3 Spectral fitting of a ULX pulsar
-In this section, we discuss the results of our fitting we obtained for the Galactic ULX pulsar Swift J0243.6+6124 (Cenko et al., 2017; Kennea et al., 2017).
-The source went into a bright outburst in 2017, with its peak luminosity $>10^{39}$ erg s-1, far exceeding the Eddington limit for a neutron star (Wilson-Hodge et al., 2018).
-We analyze the quasi-simultaneous NICER (ObsID 1050390113, 1 November 2017) and NuSTAR (ObsID 90302319004, 30 October 2017) observations that have been fitted with the standard thin-disk reflection model in Jaisawal et al.
-(2019) (see their Fig.
+TBabs $N_{H}$ $10^{22}$ cm-2 $^{*} 0.01730.0173^{*}$ $0.0173^{*}$
+zTBabs $N_{H}$ $10^{22}$ cm-2 $1.557^{+0.049}_{-0.044}$ $1.597^{+0.055}_{-0.059}$ $1.443^{+0.029}_{-0.028}$
+zcutoffpl $\Gamma$ $1.923^{+0.042}_{-0.051}$ $2.010^{+0.079}_{-0.087}$ $1.748^{+0.020}_{-0.020}$
+$E_{\rm cut}$ keV $^{*} 100100^{*}$ $100^{*}$
+norm $03440^{+0.0016}_{-0.0019} 0.0$$3680^{+0.0015}_{-0.0033} 0.0$$02874^{+0.0009}_{-0.0009}$
+reflection $h$ $M$ $5.9^{+0.9}_{-1.5}$ $2.00^{+0.003}_{-P}$
+$a$ $0.801^{+0.148}_{-0.410}$
+$\theta_{\rm half}$ deg $29.133^{+5.754}_{-4.434}$
+$v_{\infty}$ $c$ >0.384>0.384
+$R_{\rm acc}$ $M$ $10^{*}$
+$i$ deg $23.82^{P~\dagger}_{-8.71}$ $68.0^{+P}_{-1.8}$
+$\log\xi$ erg cm s-1 <0.51<0.51 $0.58^{+0.50}_{-P}$
+$A_{\rm Fe}$ solar <0.67<0.67 $0.50^{+0.08}_{-P}$
+norm $\times 10^{-5}$ $1.294^{+6.503}_{-0.385}$ $376.3^{+1160.0}_{-28.5}$
+Gaussian $E_{\rm line}$ keV $8.028^{+0.076}_{-0.078}$
+$\sigma$ keV $0.0595^{+0.127}_{-P}$
+norm $\times 10^{-5}$ $4.34^{+2.39}_{-2.02}$
+$\chi^{2}/\nu$ $108.27/124$ $111.82/125$ $143.34/128$
+Table 1: Redshift ($z$) in all models is set to $0.3534$.
+denotes the parameter was frozen and $P$ denotes the parameter reached its lower/upper limit.
+The galactic hydrogen column density ($N_{H}$) in model TBabs was obtained from https://www.swift.ac.uk/analysis/nhtot/index.php.
+The upper bound of inclination angle is always set to $\theta_{\rm half}$ in refluxlp_wind.
+The reported uncertainties correspond to the 90% interval limits for one relevant parameter ($\Delta\chi^{2}=2.71$).
+4 Discussions
+We tested our wind-reflection framework on the super-Eddington TDE Swift J1644+57.
+This source, powered by a tidal disruption of a star around a supermassive black hole, shows a highly super-Eddington accretion flow and has long been proposed as a case where powerful winds and relativistic jets coexist.
+The spectrum obtained with XMM-Newton presents clear reflection signatures in the Fe K band that cannot be adequately described by a simple Gaussian or a thin-disk reflection.
+Our analysis with refluxlp_wind suggests a wind funnel with an half-opening angle of $\sim 30^{\circ}$.
+The wind is found to be very fast, with terminal velocity $v_{\infty}\gtrsim 0.384\,c$, and the system is observed at a low inclination $i\approx 24^{\circ}$.
+The corona is compact, located at $h\sim 6~M$, which allows for effective illumination of the funnel walls while still remaining within the gravitational well of the black hole.
+These geometric and kinematic constraints together explain why the observed Fe K feature is moderately blueshifted (as can also be seen in Fig.
+5).
+The photons are scattered and reprocessed by the optically thick wind before escaping and the line-of-sight is within the funnel region.
+The opening angle and wind velocity might be related to the accretion rate as suggested from numerical simulations in Thomsen et al.
+(2019).
+We are not able to constrain the wind acceleration radius ($R_{\rm acc}$), possibly due to weak dependence of the iron line on the parameter (Fig.
+5 top-right panel) and the data quality.
+Interestingly, we notice that the wind in non-ionized ($\log\xi\ \lesssim 0.51$) and has very low iron abundance ($A_{\rm Fe}\lesssim 0.67$).
+Just for comparison and testing, we also fit the data with relxilllp and a Gaussian.
+Out of the three, the lowest $\Delta\chi^{2}$ is obtained with refluxlp_wind.
+The fitting results with relxilllp seem unphysical as the thin-disk geometry in the model predicts a high inclination of $i\sim 68^{\circ}$.
+At such a high inclination, the super-Eddington system should be obscured either by the slim-disk or the optically thick outflow.
+It also predicts extremely low corona height $h\sim 2M$ and a spin of $a\sim 0.8$.
+However, relxilllp fitting also predicts the low ionization and low iron abundance in the gas – consistent with the wind model.
+The Gaussian line fitting reveals only a narrow line ($\sigma\sim 0.06$ keV) located at $\sim 8$ keV with more residuals (Fig.
 7).
-We follow the procedure in Jaisawal et al.
-(2019) to process the NuSTAR (FPMA and FPMB) data.
-The NICER data are first processed with nicerl2 (CALDB v20240206 and geomagnetic data updated to March 2024).
-The source and background spectral files are then extracted using nicel3-spect with the default background model.
-Spectra are grouped to ensure a minimal counts of 100 for each bin before fitting with XSPEC v12.13.0c (Arnaud, 1996).
-Best-fit parameter values and uncertainties (90% confidence level unless otherwise stated) are obtained using $\chi^{2}$ statistics.
-As in Jaisawal et al.
-(2019), we first fit the NICER (1–10 keV) and NuSTAR (3–79 keV) data with the model: constant * tbabs * (bbodyrad + cutoffpl + relxill/relxilllp).
-In this model, the relxill and relxilllp (García et al., 2014) component represent relativistic reflection from a razor-thin accretion disk, assuming fixed emissivity ($q=3$) and the corona as a point source above the black hole at a certain height ($h$), respectively.
-We fix the spin parameter at $a_{*}=0$ and the outer radius of the accretion disk at 1000 $M$.
-The best-fit parameters and the corresponding uncertainties are shown in Tab.
-2 for fixed emissivity and in Tab.
-3 with lamp-post corona geometry.
-At the time of this observation, the mass accretion rate of the source was at super-Eddington level.
-The absorption corrected X-ray flux in the 0.01-100 keV band is around 10 times of the Eddington limit.
-Therefore, a thick disk, instead of thin one, is expected.
-We replace the relxill/relxilllp component with our REFLUX/REFLUXLP model, with the mass accretion rate fixed at 10 times of the Eddington accretion rate.
-As shown in Tabs.
-2 and 3, the models provide a good fit to the data, with the $\chi^{2}$ close to that of relxill/relxilllp.
-The best-fit model components and residuals are shown in Fig.
-4.
-[…]
-Tbabs nH $10^{22}$ cm-2 $0.683_{-0.011}^{+0.009}$ $0.664_{-0.006}^{+0.01}$
-Cutoffpl $\Gamma$ $1.423_{-0.009}^{+0.009}$ $1.434_{-0.009}^{+0.014}$
-Ecut keV $24.3_{-0.5}^{+0.6}$ $24.8_{-0.4}^{+0.3}$
-norm $12.4_{-0.5}^{+0.5}$ $12.40_{-0.29}^{+0.5}$
-Bbodyrad kT keV $1.213_{-0.018}^{+0.019}$ $1.234_{-0.013}^{+0.009}$
-norm $955_{-33}^{+21}$ $988_{-27}^{+25}$
-Reflection Incl deg $15.1_{-2.4}^{+4}$ $25.5_{-1.7}^{+3}$
-$R_{\rm in}$ $M$ $37_{-7}^{+9}$ $34_{-7}^{+8}$
-$\log(\xi)$ erg cm s-1 $3.54_{-0.07}^{+0.05}$ $3.29_{-0.12}^{+0.03}$
-$A_{\rm Fe}$ solar $4.2_{-1.2}^{+1.4}$ $2.39_{-0.27}^{+0.5}$
-norm $0.050_{-0.005}^{+0.007}$ $0.0255_{-0.0017}^{+0.001}$
-$C_{\rm FPMB}$ $0.9880_{-0.0014}^{+0.0014}$ $0.9880_{-0.0014}^{+0.0014}$
-$C_{\rm XTI}$ $1.242_{-0.012}^{+0.012}$ $1.240_{-0.012}^{+0.01}$
-$\chi^{2}/\nu$ 2855.14/2624 2857.12/2624
-Table 2: Best-fit parameters for the model with relxill and the model with REFLUX. The reported uncertainties correspond to the 90% interval limits for one relevant parameter ($\Delta\chi^{2}=2.71$).
-[…]
-Tbabs nH $10^{22}$ cm-2 $0.667_{-0.009}^{+0.009}$ $0.671_{-0.008}^{+0.007}$
-Cutoffpl $\Gamma$ $1.434_{-0.009}^{+0.01}$ $1.438_{-0.01}^{+0.008}$
-Ecut keV $24.74_{-0.4}^{+0.26}$ $24.7_{-0.3}^{+0.3}$
-norm $12.4_{-0.3}^{+0.4}$ $12.6_{-0.3}^{+0.3}$
-Bbodyrad kT keV $1.23_{-0.01}^{+0.01}$ $1.235_{-0.006}^{+0.005}$
-norm $1000_{-30}^{+40}$ $970_{-25}^{+30}$
-Reflection $h$ $M$ $2.20_{-P}^{+0.07}$ $>65$
-Incl deg $25.7_{-1.3}^{+2.7}$ $26.5_{-2.4}^{+1.2}$
-$R_{\rm in}$ $M$ $39_{-8}^{+9}$ $<14$
-$\log(\xi)$ erg cm s-1 $3.31_{-0.05}^{+0.04}$ $3.284_{-0.07}^{+0.029}$
-$A_{\rm Fe}$ solar $2.4_{-0.3}^{+0.3}$ $2.6_{-0.3}^{+0.4}$
-norm $3.58_{-0.7}^{+0.24}$ $1.37_{-0.06}^{+0.10}\times 10^{-7}$
-$C_{\rm FPMB}$ $0.9880_{-0.0014}^{+0.0014}$ $0.9880_{-0.0014}^{+0.0014}$
-$C_{\rm XTI}$ $1.223_{-0.013}^{+0.012}$ $1.238_{-0.01}^{+0.011}$
-$\chi^{2}/\nu$ 2816.49/2623 2826.35/2623
-Table 3: Best-fit parameters for the model with relxilllp and the model with REFLUXLP.
-The reported uncertainties correspond to the 90% interval limits for one relevant parameter ($\Delta\chi^{2}=2.71$).
-Figure 4: Best-fit model components and residuals for the lamppost geometry (left) and the power-law emissivity profile (right) with the REFLUX model.
-The black and red colors represent data of FPMA and FPMB, respectively.
-NICER data are plotted in cyan.
-Data are rebinned for visual clarity.
-4 Discussions
-The quality of the fits with the model with relxill and the model with REFLUX is similar.
-We fit the same data and the two models have the same number of free parameters, so we can directly compare their $\chi^{2}$ and the difference is marginal ($\Delta\chi^{2}=1.98$ for fixed emissivity and $9.86$ for lamppost).
-The $\chi^{2}$ of the best-fit with relxill is slightly lower – a possible reason for this could be the smaller parameter grid of REFLUX than relxill in the FITS file.
-However, we know that the source was accreting around 10 times of its Eddington limit, so the fit with REFLUX appears to be more appropriate for the spectra analyzed.
-In Tab.
-2, the two models provide consistent estimates of the parameters with the exception of the inclination angle of the disk and the iron abundance.
-The measurements of the inclination angle are $i=15.1_{-2.4}^{+4}$ deg (relxill) and $25.5_{-1.7}^{+3}$ (REFLUX).
-However, the estimates of disk inclination angles inferred from the analysis of reflection features are normally not very accurate and both models suggest a low value of $i$.
-Estimates of the iron abundance using X-ray reflection spectroscopy are even more problematic.
-It is currently unclear why reflection measurements often suggest super-solar iron abundances (but there are also sources for which reflection measurements suggest iron abundances significantly lower than the solar one); see, for instance, Bambi et al.
-(2021) for a short review on the problem.
-It may be related to a deficiency of current reflection models (e.g., underestimate of the disk electron density (Jiang et al., 2019a, b; Liu et al., 2023b), presence of two or more coronae (Fürst et al., 2015), etc.) or a true physical phenomenon, like radiative levitation of metal ions in the inner part of the accretion disk (Reynolds et al., 2012).
-The estimate of the iron abundance with REFLUX would go to the right direction, decreasing the value of $A_{\rm Fe}$ with respect to that inferred with relxill, but we would need to analyze more sources and more observations to see if this is a feature of our model and not an accidental result for the spectra analyzed in the previous section.
-In Tab.
-3, we see that for a lamppost geometry, the inclination angle and the iron abundance for both the models closely match the result of REFLUX in the analysis with fixed emissivity.
-However, we see a significant difference in the corona height and the inner edge of the disk.
-With relxill, we see a very low corona height (the lowest in the model) and a larger value for the inner edge.
-With REFLUX, the case is the opposite, we obtain a higher corona height $h>65M$ and inner edge closer to the central object $R_{in}<14M$.
-In the case of REFLUX, we see a small norm possibly because the whole disk may not be illuminated for many different $h$ and $R_{in}$ configurations making the flux from the outer part of the disk negligible.
-We caution the readers against interpreting the high value of $h>65M$ too rigidly.
-As in our analysis, we assume a lamppost geometry, which may not be applicable to an Ultra-Luminous Pulsar (ULP) like Swift J0243.6+6124.
-The system might instead contain a more extended primary illuminating source, such as an accretion column (e.g., Pinto & Walton (2023)).
-Nonetheless, we observed changes in inferred geometries when a slim disk model was employed.
-This underscores the critical importance of considering disk geometry when addressing systematic uncertainties.
-Such uncertainties significantly affect the inferred disk inner radius and, consequently, the magnetic field of ULPs (e.g., Jaisawal et al.
-(2019)).
-We will further investigate Swift J0243.6+6124 in this perspective in the future.
-A more detailed analysis of the 2017 NICER and NuSTAR data of Swift J0243.6+6124 is beyond the purpose of the present work, which is the presentation of REFLUX.
-In our fits (with both relxill and REFLUX), we have assumed that the spectrum illuminating the disk, and responsible for the reflection component can be described by a power law with a high-energy cutoff.
-However, this may not be the case for Swift J0243.6+6124.
-A power law with a high-energy cutoff is the spectrum expected for Comptonized photons, which is the case in which thermal photons from an accretion disk can inverse Compton scatter off free electrons in a hot corona.
-In the case of a neutron star, reflection features may be produced by illumination of the disk by thermal photons from the surface of the neutron star.
-Such a scenario could be implemented in REFLUX by using the xillverNS table, in which reflection spectra are produced by an incident blackbody spectrum.
-While for neutron star XRBs like Swift J0243.6+6124 the Schwarzschild metric should describe well the spacetime geometry around the source, in the case of other super-Eddington X-ray sources (black hole XRBs, AGNs, and TDEs) deviations from the Schwarzschild background can be important and it will be necessary to implement the Kerr metric in REFLUX.
-This is also left to future work.
-This work was supported by the National Natural Science Foundation of China (NSFC), Grant No.
-12250610185 and 12261131497, and the Natural Science Foundation of Shanghai, Grant No.
-22ZR1403400.
-J.J.
-acknowledges support from Leverhulme Trust, Isaac Newton Trust and St Edmund’s College, University of Cambridge.
+The fitting results of zTBabs and zcutoffpl are consistent across the three fitting results, with only the fit with zGauss having slight discrepancy (see Tab.
+1).
+From an astrophysical perspective, the wind geometry inferred for Swift J1644+57 is consistent with hydrodynamic and magnetohydrodynamic (MHD) simulations of super-Eddington accretion, which predict fast, optically thick outflows launched from the inner disk.
+The funnel angle and velocity we measure are comparable to those suggested by radiation-MHD models.
+A misclassification of wind reflection as disk reflection would bias spin or inclination estimates, highlighting the necessity of using models that capture the relevant geometry.
+A more detailed analysis of the 2011 XMM-Newton data of Swift J1644+57 is beyond the purpose of the present work, which is the presentation of reflux.
+The wind-reflection fit of Swift J1644+57 emphasizes that super-Eddington systems can show fundamentally different reflection morphologies depending on whether the inner flow geometry is dominated by a geometrically thick disk or by a massive outflow.
+In all our fits, we have assumed that the spectrum illuminating the disk and responsible for the reflection component can be described by a power law with a high-energy cutoff.
+However, in the case of a TDE the reflection might be produced from the returning radiation of the thermal component from the disk, in such a case using xillverTDE table might be useful (Yao et al., 2024) but even better would be to have a reflection model which accounts for full returning radiation of thermal emissions (e.g.
+Mirzaev et al., 2024b, but with a super-Eddington geometry).
+Nevertheless, the TDE Swift J1644+57 was observed to have strong jet (Burrows et al., 2011), in such a case the base of the jet could act as the corona and our assumption stands well.
+Our work motivates a broader application of both the slim-disk and wind-reflection models to a wide range of super-Eddington sources.
+These models applied with other spectral, timing and polarimetric data could really help decode the nature of accreting matter in extreme regimes.
+In future work, we plan to analyze data of sources which could possibly have the accretion geometry as in reflux.
+In terms of modeling, we wish to update reflux with the Kerr metric opening up the possibility of measuring black hole spins.
+We also hope to update the geometries with precise numerical accretion models.
+We would like to thank the anonymous reviewer for their constructive comments to improve the manuscript content.
+This work was supported by the National Natural Science Foundation of China (NSFC), Grant Nos.
+12250610185, W2531002, and 12261131497.
+$S.$ acknowledges support from Shanghai Super Postdoctoral fellowship.
+$J.$ acknowledges support from Leverhulme Trust, Isaac Newton Trust and St Edmund’s College, University of Cambridge.
+Part of the computations in this research were performed using the CFFF platform of Fudan University.